what is the parametric (point) form of the parabola x^2=4ay.

Grade:11

2 Answers

shivam kumar
7 years ago
Parametric form of parabola y^2=4ax are x=at^2 , y=2at where t is a parameter at any point of this parabola can be taken as P=(at^2,2at).
shivam kumar
7 years ago
Parametric equation of the parabola x^2=4ay are x=2at and y=at^2 where t is a parameter at any point of this parabola can be taken as P=(2at,at^2).